OpenStudy (anonymous):

The Big Bull investment club has a portfolio that contains only two stocks. At the end of the year, stock A had gone up 8%, and stock B had gone up 10%. The total portfolio, which had a value of $10,000 at the beginning of the year, went up 9.2%. How much is the investment in stock A? @ßαbчღ

OpenStudy (anonymous):

.Equation: interest + interest = interest 0.08x + 0.10(10,000-x) = 0.092(10,000) Multiply thru by 1000 80x + 100*10,000-100x = 92*10,000 -20x = -8*10,000 x = $4000 (amt invested at 8%) --- 10,000-x = $6000 (amt invested at 10%

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Welcome

OpenStudy (anonymous):

i am ganna posted another and i alreday ganna take u in